# SFFT-based Homogenization: Using Tensor Trains to Enhance FFT-Based Homogenization

### v1

@@ 5.3 Solution Fields
The stress solution fields obtained using the SFFT-based Homogenization algorithm were compared to the FFT-based approach for both the Box and Circle geometries, see Figure 10). For both geometries, the relative error in the solution fields is greatest at the boundaries of the inclusion. Interestingly, for the Circle geometry, rectangular regions with similar relative error are observed. This suggests that the cQTT format is better suited for handling straight, non-curved edges in the geometry. To accommodate curved geometries within the SFFT-based Homogenization framework, alternative TN architectures, such as PEPS, MERA, or other generalizations of TTNs, may prove beneficial. However, this analysis lies outside the scope of the current paper and will be addressed in future work.
@@ 6 Conclusion
In this paper, we introduce the SFFT-based Homogenization, a geometry-tailored adaptation of the FFT-based Homogenization method that leverages the cQTT variant of the QFT, known as the SFFT. This novel algorithm offers key advantages, including improved time complexity and reduced memory consumption, outperforming the traditional FFT-based approach in scenarios where the geometry can be effectively compressed into a low-rank cQTT format. The advantage stems from the tailored nature of our method in contrast to the ’one-size-fits-all’ approach of the FFT-based method. Notably, memory consumption no longer scales exponentially and consistently outperforms the FFT-based approach in relevant cases.
While the FFT-based approach retains an advantage for smaller discretizations, the SFFT-based algorithm surpasses it beyond a geometry-dependent crossover. Beyond this point, the improved time complexity of our method leads to a substantial speedup. These benefits are particularly pronounced in 3D settings with compressible 2D or 1D structures, where the additional dimensions represent redundant data that the SFFT can efficiently handle.
However, the current version of our algorithm struggles with geometries featuring circular inclusions, possibly due to the specific choice of the underlying TN architecture. This raises the question of whether alternative network architectures, such as PEPS or TTN, could address this shortcoming.
Leveraging GPUs for high discretizations, made possible by reduced memory consumption, presents another potential enhancement [35]. This capability is infeasible for traditional FFT-based Homogenization at high discretizations. Similarly, utilizing TPUs [36] and optimizing the underlying software could provide substantial speedups, potentially offering a decisive advantage to the SFFT-based Homogenization algorithm. This is further supported by the observation that the anticipated bottleneck imposed by the SFFT has not yet been reached, with runtime instead dominated by the multiplication of the Green-Eshelby operator and the intermediate strain.
@@ Acknowledgments
We thank BMWK for the financial support provided under the EniQmA project. Furthermore, we acknowledge the funding of the German National High Performance Computing (NHR) Association for the Center NHR South-West.
@@ Appendix A Greens operator Discretization Schemes
The discretization of computational domains and associated variables is a critical step in numerical modelling, particularly in Fourier-based methods. This appendix provides a detailed discussion of the Fourier-space discretization schemes relevant to FFT- and SFFT-based Homogenization. These schemes play a central role in ensuring accurate and efficient computations, particularly by addressing numerical artifacts and ensuring physical consistency in derived quantities.
@@
$\varepsilon_{ij}=(\mathbf{q})\frac{1}{2}\left[k_{i}(\mathbf{q})u_{j}(\mathbf{q})+k_{j}(\mathbf{q})u_{i}(\mathbf{q})\right].$
@@
In continuous space, the Green-Eshelby operator exhibits the symmetry $\hat{\varGamma}^{0}(\mathbf{q})^{\ast}=\hat{\varGamma}^{0}(\mathbf{-q})$, which ensures that the physically observable strain $\varepsilon$ is real-valued. For an odd number of discretization points, this symmetry is inherently preserved in the discretized version of the operator. However, for an even number of discretization points, the Nyquist frequency becomes part of the spectrum but lacks a negative counterpart. This may result in symmetry breaking, which in this context is a purely numerical artifact that must be addressed accordingly. The approach we follow enforces symmetry at the Nyquist frequencies $q_{ny}$ explicitly through
@@ Appendix B Approximate Tensor Train Arithmetic
A significant class of TT-TTO operations consist of optimization algorithms to solve systems of linear equations or to obtain an approximate solution to a matrix-vector product. There are many different approaches that were developed to solve these problems. The most famous ones are the alternating least square (ALS) method [37], the density matrix renormalization group (DMRG) algorithm [38] and the alternating minimal energy (AMEn) algorithm [30]. The latter one being regarded as the current state-of-the-art method.
However, instead of using the AMEn algorithm to solve liner systems of equations, we will use an adapted method for faster matrix-vector products.
@@
It can readily be verified that the gradient of the cost function is zero when the following corresponding Galerkin condition is met
@@
In the third step, the residual $z=\tilde{y}-Ax$ is computed. This step is generally computationally expensive and thus undesirable; therefore, an approximate residual $\tilde{z}\approx z$ is used instead. Once the approximate residual is obtained, its first core is used to enrich the original first core of the solution $y_{1}$. This step, known as the Galerkin correction, not only allows for incremental refinement but, more importantly, enables an adaptive rank procedure.
It is important to note that this step is not exclusive to the AMEn algorithm; similar approaches can be found in other solvers as well.
@@
The main advantage of this class of optimization methods, compared to an exact solution, lies not only in its improved time complexity but also in the resulting rank. If the result $y$ has a rank similar to the input TT $x$, the time complexity improves by a factor of $r(A)^{2}$, Moreover, the rounding procedure becomes noticeably faster, as it depends on the cubic rank. With a quasi-exact solution, the final rank scales with the product of the matrix and vector ranks.
We conclude this section by noting that we have used a revised version of the TTO-TT matrix-vector product from the TT-Toolbox available on GitHub.
@@ Appendix D Scalable Implementation
The implementation of the SFFT algorithm needed to achieve a certain scalability threshold for high-resolution scenarios. In this case, characterized by a large number of discretization points per dimension, we observed significant memory and time consumption during the SFFT-based algorithm’s preparation phase. This issue arises due to the reliance on the state-of-the-art TT-SVD algorithm, which is constrained by the computational bottleneck of the SVD.
@@
As such, the SVD step scales with a complexity of $\mathbf{O}\left(mn^{2}\right)$ for an $m\times n$ matrix, where $n<m$. Since higher-resolution matrices scale exponentially in size, this results in exponential time consumption for separating the cores in the TT or TTO representations. Such behaviour is highly undesirable, as it merely shifts the computational burden from the iterative phase of the algorithm to the preprocessing phase.
Thus, instead of using the TT-SVD method, we will use a combination of the streamable tensor-train approximation (STTA) [39] and an updated TT-SVD method build on the randomized SVD. The specific choice of the above methods depends on the computational and memory restrictions of the compute system used.
@@
The STTA approach can be used in a multi-core setting using pythons dask library [40] and is especially useful for matrices with a lower and intermediate resolution. In the high-resolution regime, we are typically encountering to high memory restrictions for the multi-core system to be beneficial. This is either due to the system not having enough memory per core or due to the high overhead of transferring the compute-graph between the nodes. In this case, the slower non-multicore version of the randomized TT-SVD with additional rank truncation threshold can be used. While the SVD of a rank $k$ matrix with shape $mxn$ has a time-complexity $\mathcal{O}\left(mn*\min(m,n)\right)$ that scales quadratic, the randomized SVD only scales linear in the matrix dimensions according to $\mathcal{O}\left(mn*\log(k)+\left(m+n\right)*k^{2}\right)$ [41]. Thus, we are able to speed up the process by at least a quadratic factor, which represents a significant improvement for large matrices. Alternatively, the TT-cross approximation can be used. The TT-cross method relies on a cross-approximation strategy, where the most important components of the tensor are iteratively selected to provide the best rank-reduction approximation. TT-cross selects a small subset of the tensor’s entries for its approximation, which efficiently captures the tensors structure [34]. However, since (randomized) TT-SVD is more accurate, we opted not to use TT-cross for obtaining the optimal solutions for our derived TT.
Additionally, we need to consider the case of reusing preprocessed computations for subsequent runs.
@@
Almost all TTs and TTOs within the SFFT-based homogenization algorithm depend only on the dimensionality, the chosen resolution, and the core ordering. This includes both the (inverse) SFFT and the initial strain TT. As a result, these components can be saved in a lookup table for later use across multiple problems that share a common set of system parameters. A special case in this context is the geometry-dependent stiffness operator $C(x,y)$ as well the Green-Eshelby operator $\hat{\Gamma}_{0}(k_{x},k_{y})$. The latter depends on the chosen Lamé parameters as well as the discretization scheme used. Since the Green-Eshelby operator inherently defines the discretization choice for the problem, recomputation is generally unavoidable for different schemes. However, this is not necessary for varying sets of Lamé parameters $\{\mu,\lambda\}$. In such cases, we can reuse the Green-Eshelby operator, which can generally be expressed in the following form
$\hat{\Gamma}_{0}(k_{x},k_{y})=\sum_{i}\alpha_{i}(\mu,\lambda)\hat{\Gamma}^{(i)}(k_{x},k_{y}).$ (18)
@@
Therefore, the reduced Gamma functions $\hat{\Gamma}^{(i)}(k_{x},k_{y})$ can be stored separately in TT format for efficient reuse. When needed, the prefactors $\alpha_{i}(\mu,\lambda)$ can easily be computed and the saved reduced Gamma functions can be recombined according to equation (18) into the full form through simple TT arithmetic can be easily computed, and the stored reduced Gamma functions can be recombined according to equation (18) into the full form using simple TT arithmetic.
@@
Finally, in contrast to all other quantities in the SFFT-based Homogenization algorithm, the stiffness operator $C(x,y)$ must be recomputed each time a new geometry is used.
Appendix E Additional Geometries
@@
Figure 11: Fundamental 2D geometries used to probe the regimes of the SFFT-based Homogenization algorithm: a) Checkerboard, b) Spots and c) Voronoi. The grey regions have a Young’s Modulus of $E_{1}=29/3\,$$\mathrm{GPa}$ while the white regions display $E_{2}=4/3\,$$\mathrm{GPa}$. The Poisson ratios were set to $\nu_{1}=\nu_{2}=1/3$ for the grey and white regions, respectively.
@@
Figure 14: Maximal Strain Rank $r_{\mathrm{s}}$ during the evaluation of the algorithm in dependence of the discretization order $n$ in a) 2D and b) 3D space
@@
For the Voronoi geometry, a large number of points was initially sampled, and the Voronoi regions for each point were constructed. The final inclusions were obtained by randomly selecting a fraction of these points, in accordance with the predefined volume fraction $\Phi_{vol}=0.9$. The material parameters for the selected Voronoi regions were then assigned the inclusion parameters, as shown in Figure 11. For more detailed information on the construction of the Voronoi geometry, refer to [42].
@@
The memory usage results, presented in Figure 12, demonstrate the stark contrast between the exponential scaling of the FFT-based method and the geometry-optimized efficiency of the SFFT approach in both 2D and 3D. For the Checkerboard and reduced Checkerboard geometries, the SFFT method significantly reduces memory consumption. On the other hand, the Voronoi and Voronoi 3D geometries exhibit exponential memory growth, similar to the FFT method. For generalized geometries like the Checkerboard 3D, the SFFT method still offers superior memory efficiency compared to the FFT baseline. More complex geometries like the Spots, Voronoi, Fibres or Voronoi 3D Geometries all show exponential memory scaling similar to the state-of-the-art FFT based Method. However, due to the redundant z-component of the Fibres geometry, we still see a considerable memory reduction in comparison to the FFT-based method.
@@
Regarding iteration times, as shown in Figure 13, we observe that the FFT method is faster for lower discretization orders. However, for the from the Checkerboard derived geometries, the SFFT method starts to outperform FFT after reaching a geometry-dependent threshold. In 2D, this crossover occurs at $n=15$ for the Checkerboard geometry, while in 3D, the reduced Checkerboard geometry shows a crossover around $n=10$. The cross over point for the Checkerboard 3D geometry is expected to be araound $n=15$.
@@
In Figure 14, we see that the rank growth of the Checkerboard and all its two generalizations into 3D space exhibit sub-exponential scaling, while Voronoi, Spots and their 3D counterpart all exhibit exponential rank scaling.
@@
Figure 15 illustrates the dependence of relative error on the discretization order $n$. Here, all geometries reach accuracies of $10^{-5}$ or better.
@@
In conclusion, these additional geometries further emphasize that the current SFFT-based Homogenization algorithm is particularly well-suited for rectangular geometries. However, the results also highlight that geometries incompatible with the existing tensor network structure exhibit scalability performance similar to that of the traditional FFT-based method.
Appendix F Discretization Scheme

### v3

@@ 5.3 Solution Fields
The stress solution fields obtained using the SFFT-based Homogenization algorithm were compared to the FFT-based approach for both the Box and Circle geometries, see Figure 10. For both shapes, the relative error in the solution fields is greatest within the inclusion. Interestingly, for the Circle geometry, a multitude of rectangular regions with similar relative error are observed. This suggests that the TT format is better suited for handling straight, non-curved edges in the geometry. To accommodate curved geometries within the SFFT-based framework, alternative Tensor Network architectures, such as projected entangled pair states (PEPS), the multi-scale entanglement renormalization ansatz (MERA), or other generalizations of Tree Tensor Networks [46, 47, 48], can be considered. This approach may be advantageous, as different Tensor Network formats encode the same structured data in different ways, leading to varying compression efficiencies [49]. However, this analysis is beyond the scope of the present paper and will be explored in future work.
@@ 6 Conclusion
In this paper, we introduce the SFFT-based Homogenization algorithm, a geometry-tailored adaptation of the FFT-based Homogenization method that leverages the TT variant of the QFT, known as the SFFT. This novel algorithm offers key advantages, including improved time complexity and reduced memory consumption, outperforming the FFT-based approach in scenarios where the geometry can be effectively compressed into a low-rank TT format. The advantage stems from the tailored nature of our method, in contrast to the ’one-size-fits-all’ approach of the FFT-based algorithm. In particular, memory consumption no longer scales exponentially and consistently outperforms the FFT-based Homogenization algorithm in relevant cases.
While the FFT-based approach retains an advantage for smaller discretizations, the SFFT-based algorithm surpasses it beyond a geometry-dependent crossover. Beyond this point, the improved time complexity of our method leads to a substantial speedup. These benefits are particularly pronounced in 3D settings with compressible 2D or 1D structures, where the additional dimensions represent redundant data that is processed with high efficiency. Additionally, the derived time complexity of our algorithm $\mathcal{O}\Big{\{}nD\,r_{\mathrm{S}}^{3}\left(r_{\mathrm{C}}^{3}+r_{\varGamma}\right)\Big{\}}$, further reinforces the potential for exponential speedup, provided the contributing ranks remain small.
However, the current version of our algorithm struggles with geometries featuring circular inclusions, possibly due to the specific choice of the underlying Tensor Network architecture. This raises the question of whether alternative network architectures, such as PEPS, MERA or other Tree Tensor Networks, could address this shortcoming.
Leveraging GPUs for high discretizations, made possible by reduced memory consumption, presents another potential enhancement [50]. This capability is infeasible for traditional FFT-based Homogenization at high discretizations. Similarly, utilizing TPUs [51] and optimizing the underlying software could provide substantial speedups, potentially offering a decisive advantage to the SFFT-based Homogenization algorithm. This is further supported by the observation that the current bottleneck of our algorithm is not due to the theoretical limits of the SFFT, but rather the optimization method applied to the matrix-vector product between the Green-Eshelby operator $\hat{\varGamma}_{0}$ and the polarization iterates $\hat{\tau}_{m}$. Finally, we emphasize that our algorithm’s results have the potential to significantly accelerate data generation for deep material networks, enabling the prediction of nonlinear elastic material properties from linear elastic data.
@@ Acknowledgments
To enhance readability and ensure comprehensiveness, portions of this manuscript were refined with the assistance of FhGenie. We thank BMWK for the financial support provided under the EniQmA project. Furthermore, we acknowledge the funding of the German National High Performance Computing (NHR) Association for the Center NHR South-West.
@@ Appendix A Greens operator Discretization Schemes
The discretization of computational domains and associated variables is a critical step in numerical modelling, particularly in Fourier-based methods. This appendix provides a detailed discussion of the Fourier-space discretization schemes relevant to FFT- and SFFT-based Homogenization algorithms. These schemes play a central role in ensuring accurate and efficient computations, particularly by addressing numerical artifacts and ensuring physical consistency in derived quantities.
@@
$\varepsilon_{ij}(\mathbf{q})=\frac{1}{2}\left[k_{i}(\mathbf{q})u_{j}(\mathbf{q})+k_{j}(\mathbf{q})u_{i}(\mathbf{q})\right].$
@@
In continuous space, the Green-Eshelby operator exhibits the symmetry $\hat{\varGamma}^{0}(\mathbf{q})^{\ast}=\hat{\varGamma}^{0}(\mathbf{-q})$, which ensures that the physically observable strain $\varepsilon$ is real-valued. For an odd number of discretization points, this symmetry is inherently preserved in the discretized version of the operator. However, for an even number of discretization points, the Nyquist frequency becomes part of the spectrum but lacks a negative counterpart. This results in symmetry breaking, which is a purely numerical artifact that must be addressed accordingly. The approach we follow enforces symmetry at the Nyquist frequencies $\mathbf{q}_{ny}$ explicitly through
@@ Appendix B Approximate Tensor Train Arithmetic
A significant class between TTs and TTOs consist of optimization algorithms to solve systems of linear equations or to obtain an approximate solution to a matrix-vector product. There are many different approaches that were developed to solve these problems. The most famous ones are the alternating least square (ALS) method [52], the density matrix renormalization group (DMRG) algorithm [53] and the alternating minimal energy (AMEn) algorithm [38]. The latter one being regarded as the current state-of-the-art method. Here we will use an adaptation of the classical AMEn algorithm to solve the matrix-vector product of a TT with a TTO.
@@
It can readily be verified that the gradient of the cost function is zero when the corresponding Galerkin condition is met
@@
In the third step, the residual $z=\tilde{y}-Ax$ is computed. This step is generally computationally expensive and thus undesirable; therefore, an approximate residual $\tilde{z}\approx z$ is used instead. Once the approximate residual is obtained, its first core is used to enrich the original first core of the solution $y_{1}$. This step, known as the Galerkin correction, not only allows for incremental refinement but, more importantly, enables an adaptive rank procedure. It is important to note that rank adaptation is not exclusive to the AMEn algorithm; similar approaches can be found in other solvers as well.
@@
The main advantage of this class of optimization methods, compared to an exact solution, lies not only in its improved time complexity but also in the resulting rank. If the result $y$ has a rank similar to the input TT $x$, the time complexity improves by a factor of $r(A)^{2}$, Moreover, the rounding procedure becomes noticeably faster, as it depends on the cubic rank. With a quasi-exact solution, the final rank scales with the product of the matrix and vector ranks. Finally, we want to note that we have used an improved version of the TTO-TT matrix-vector product built on the matrix-vector within the TT-Toolbox available on GitHub.
@@ Appendix D Scalable Implementation
The implementation of the SFFT-based Homogenization algorithm needed to achieve a certain scalability threshold for high-resolution scenarios. In this case, characterized by a large number of discretization points per dimension, we observed significant memory and time consumption during the SFFT-based algorithm’s preparation phase. This issue arises due to the reliance on the state-of-the-art TT-SVD algorithm [37], which is computationally constrained by the usage of the SVD.
@@
As such, the SVD step scales with a complexity of $\mathbf{O}\left(mn^{2}\right)$ for an $m\times n$ matrix, where $n<m$. Since higher-resolution matrices scale exponentially in size, this results in exponential time consumption for separating the cores in the TT or TTO representations. Such behaviour is highly undesirable, as it merely shifts the computational burden from the iterative phase to the preprocessing phase. Thus, instead of using the TT-SVD method, we will use a combination of the streamable tensor-train approximation (STTA) [54] and an updated TT-SVD method build on the randomized SVD. The choice between these methods depends on the computational and memory restrictions of the compute system used.
@@
The STTA approach can be used in a multi-core setting using pythons dask library [55] and is especially useful for matrices with a lower and intermediate resolution. In the high-resolution regime, we are typically encountering too high memory restrictions for the multi-core system to be beneficial. This is either due to the system not having enough memory per core or due to the high overhead of transferring the compute-graph between the nodes. In this case, the slower non-multicore version of the randomized TT-SVD with additional rank truncation threshold can be used. While the SVD of a rank $k$ matrix with shape $m\times n$ has a time complexity $\mathcal{O}\left(mn*\min(m,n)\right)$ that scales quadratic, the randomized SVD only scales linear in the matrix dimensions according to $\mathcal{O}\left(mn*\log(k)+\left(m+n\right)*k^{2}\right)$ [56]. Thus, we are able to speed up the process by at least a quadratic factor, which represents a significant improvement for large matrices. Alternatively, the TT-cross approximation can be used. In this case, a cross-approximation strategy is employed, where the most important components of the tensor are iteratively selected to provide the best reduced rank approximation. TT-cross selects a small subset of the tensor’s entries for its approximation, which efficiently captures the tensors structure [45]. However, since (randomized) TT-SVD is more accurate, we opted not to use TT-cross for obtaining the optimal solutions for our derived TT.
@@
To improve scalability, we need to consider the case of reusing preprocessed computations for subsequent runs. Almost all TTs and TTOs within the SFFT-based Homogenization algorithm depend only on the dimensionality, the chosen resolution, and the core ordering. This includes both the (inverse) SFFT and the initial strain TT. As a result, these components can be saved in a lookup table for later use across multiple problems that share a common set of system parameters. A special case in this context is the geometry-dependent stiffness operator $\mathbf{C}(\mathbf{x})$ as well the Green-Eshelby operator $\hat{\varGamma}_{0}(\mathbf{k})$. The latter depends on the chosen Lamé parameters as well as the discretization scheme used. Since the Green-Eshelby operator inherently defines the discretization choice for the problem, recomputation is generally unavoidable for different schemes. However, this is not necessary for varying sets of Lamé parameters $\{\mu,\lambda\}$. In such cases, we can reuse the Green-Eshelby operator, which can generally be expressed in the following form
$\hat{\varGamma}_{0}(\mathbf{k})=\sum_{i}\alpha_{i}(\mu,\lambda)\hat{\varGamma}^{(i)}(\mathbf{k}).$ (18)
@@
Therefore, the reduced Gamma functions $\hat{\varGamma}^{(i)}(\mathbf{k})$ can be stored separately in TT format for efficient reuse. When needed, the prefactors $\alpha_{i}(\mu,\lambda)$ can easily be computed and the saved reduced Gamma functions can be recombined according to equation (18).
@@
Finally, in contrast to all other quantities in the SFFT-based Homogenization algorithm, the stiffness operator $\mathbf{C}(\mathbf{x})$ must be recomputed each time a new geometry is used.
Appendix E Time Complexity
For the following discussion, we distinguish between the following ranks
  • $r_{\mathrm{S}}$ for the strain TT,
  • $r_{\mathrm{\tau}}$ for the strain TT in momentum space,
  • $r_{\mathrm{C}}$ for the raised reduced stiffness TTO $\left[\mathbf{C}-\mathbf{C}(\mathbf{x})\right]$,
  • $r_{\mathrm{F}}^{(i)}$ for the (inverse) SFFT TTO $\mathcal{S}^{(-1)}$,
  • $r_{\varGamma}$ for the raised Green-Eshelby Operator TTO $\hat{\varGamma}_{0}$.
While $r_{\mathrm{F}}^{(i)}$ is determined solely by the chosen resolution, $r_{\mathrm{\varGamma}}$ also depends on the discretization scheme used. The rank $r_{\mathrm{C}}$ is geometry dependent and grows with increasing structural complexity. The strain rank $r_{\mathrm{S}}$ is influenced by all previously mentioned factors. We assume that during the initial few iterations, $r_{\mathrm{S}}$ grows until it reaches an approximate saturation point, after which it remains stable in the spatial domain. The rank can still change slightly during the iterative process, however, these changes are negligibly small. In momentum space, we make a similar assumption, but instead of matching the spatial rank, we assume it follows a linear relation $r_{\mathrm{\tau}}=ar_{\mathrm{S}}$, with some small transfer constant $a$. Finally, the mean-strain TT exhibits a trivial rank of 1 and is thus negligible. Using these assumptions on algorithm 2 and the data from Table 2, we can readily derive the time complexity of our novel SFFT-based Homogenization algorithm as
$\mathcal{O}\Big{\{}nD\,r_{\mathrm{S}}^{3}\left[\left(1+a^{3}\right)r_{\mathrm{f}}^{3}+r_{\mathrm{C}}^{3}+ar_{\varGamma}\right]\Big{\}}.$ (19)
We assume that the SFFT and its inverse have approximately the same rank, as validated in Section 5. We also use the geometric dimensionality $D$ instead of the core dimensionality $d$, since the largest core dimension determines the time complexity. In our quantized TT approach, the first core, governed by the Voigt formalism, sets the largest core dimension. For the small $D$ regime, this leads to an approximately linear relationship between $D$ and the Voigt dimensionality of the first core, with an additional multiplicative offset. The dominant factor in computational complexity is the strain rank $r_{\mathrm{S}}$, influenced by the solution field smoothness and geometry. We simplify the time complexity derived in (19) using empirical data from Section 5. First, the transfer constant satisfies $a\approx 1$. Second, the significantly lower rank of the (inverse) SFFT TTO compared to the strain-field and Green-Eshelby operator shows that the SFFT contribution is negligible small. This results in the simplified time complexity of algorithm 2 given by
$\mathcal{O}\Big{\{}nD\,r_{\mathrm{S}}^{3}\left(r_{\mathrm{C}}^{3}+r_{\varGamma}\right)\Big{\}}.$ (20)
The above reduced time complexity shows a significant dependency on the geometry since both the ranks $r_{\mathrm{S}}$ and $r_{\mathrm{C}}$ depend on it. Furthermore, since both ranks have a cubic contribution, it indicates that only for geometries with low rank the asymptotic scaling of our novel algorithm is beneficial.
@@ Appendix E Time Complexity
Figure 11: Additional 2D geometries used to probe the regimes of the SFFT-based Homogenization algorithm: a) Checkerboard, b) Spots and c) Voronoi. The grey regions have a Young’s Modulus of $E_{1}=29/3\,$$\mathrm{GPa}$ while the white regions display $E_{2}=4/3\,$$\mathrm{GPa}$. The Poisson ratios were set to $\nu_{1}=\nu_{2}=1/3$ for the grey and white regions, respectively.
Appendix F Additional Geometries
@@
Figure 14: Maximal Strain Rank $r_{\mathrm{S}}$ during the evaluation of the algorithm in dependence of the discretization order $n$ in a) 2D and b) 3D space.
@@
For the Voronoi geometry, a large set of points was initially sampled, followed by the construction of their corresponding Voronoi regions. The final inclusions were obtained by randomly selecting a fraction of these points, in accordance with the predefined volume fraction $\Phi_{vol}=0.9$. The material parameters for the selected Voronoi regions were then assigned to the inclusion parameters, as shown in Figure 11. For a more detailed derivation on the construction of the Voronoi geometry, we refer the interested reader to [57].
@@
The memory usage results, presented in Figure 12, demonstrate the stark contrast between the exponential scaling of the FFT-based method and the geometry-optimized efficiency of the SFFT-based approach in both 2D and 3D. For the Checkerboard and reduced Checkerboard geometries, the SFFT method significantly reduces memory consumption. On the other hand, the Voronoi and Voronoi 3D geometries exhibit exponential memory growth, similar to the FFT-based method. For generalized geometries like the Checkerboard 3D, the SFFT method still offers superior memory efficiency compared to the FFT baseline. More complex geometries like the Spots, Voronoi, Fibres or Voronoi 3D geometries all show exponential memory scaling similar to the state-of-the-art FFT based Method. However, due to the redundant z-component of the Fibres geometry, we still see a considerable memory reduction in comparison to the FFT-based method.
@@
For the iteration times, as shown in Figure 13, we observe that the FFT-based method is faster for lower discretization orders. However, for all the from the Checkerboard derived geometries, the SFFT-based method starts to outperform FFT-based Homogenization algorithm after reaching a geometry-dependent crossover point. In 2D, the crossover occurs at $n=15$ for the Checkerboard geometry, while in 3D, the reduced Checkerboard geometry shows a crossover around $n=10$. For the full 3D Checkerboard geometry, the crossover is expected near $n=15$, however, this extrapolation extends too far beyond the tested range to be considered reliable.
@@
In Figure 14, we see that the rank growth of the strain TT for the Checkerboard and its two generalizations exhibit sub-exponential scaling, while Voronoi, Spots and their 3D counterparts exhibit exponential rank growth. This is consistent with the previous results on iteration time and memory consumption.
@@
The dependants of the relative error on the discretization order $n$ can be seen in Figure 15. Here, all geometries reach accuracies of $10^{-5}$ or better. Thus, they follow the same trend as the originally considered geometries in Section 5.
@@
In conclusion, these additional geometries further emphasize that the current SFFT-based Homogenization algorithm is particularly well-suited for rectangular geometries. However, the results also highlight that geometries incompatible with the existing tensor network structure exhibit scalability performance similar to that of the traditional FFT-based method or worse.
Appendix G Discretization Scheme